Presented below are selected transactions at Ridge Company for 2017.

Jan. 1 Retired a piece of machinery that was purchased on January 1, 2007. The machine cost $63,500 on that date. It had a useful life of 10 years with no salvage value.
June 30 Sold a computer that was purchased on January 1, 2014. The computer cost $39,300. It had a useful life of 5 years with no salvage value. The computer was sold for $15,400.
Dec. 31 Discarded a delivery truck that was purchased on January 1, 2013. The truck cost $42,900. It was depreciated based on a 6-year useful life with a $3,000 salvage value.


Journalize all entries required on the above dates, including entries to update depreciation, where applicable, on assets disposed of. Ridge Company uses straight-line depreciation. (Assume depreciation is up to date as of December 31, 2016.)
Answer #1

Ans: Journal Entries in thr books of Ridge Company for 2017

1. Retired a piece of machinery that was purchased on January 1, 2007. The machine cost $63,500 on that date. It had a useful life of 10 years with no salvage value.

No proceeds, Asset fully depreciated and no salvage value.

Date Account Name Debit ($) Credit ($)
Jan 1, 2017 Accumulated depreciation (see working) $63500
Machinery $63500
(Being Machinery disposed off)

Working:

Depreciation per year= Cost of the asset - Salvage Value / Life of the asset = $63500/10 = $6350

Depreciation from Jan 1. 2007 to Jan 1, 2017 (i.e 10 years) = $6350 * 10 = $63500

2. Sold a computer purchased on January 1, 2014. The computer cost $39,300. It had a useful life of 5 years with no salvage value. The computer was sold for $15,400.

Date Account Name Debit ($) Credit ($)
June 30,2017 Cash $15400
Accumulated Depreciation $27510
Gain on asset disposal (Working) $3610
Computer $39300
(Being gain recored on disposal of computer)

Working:

Depreciation on computer from (Jan 1, 2014 to June 30, 2017) i.e 3.5 years = Cost of the asset - Salvage Value / Life of the asset = $39300/5= $7860 * 3.5 years = $27510

Written down value of the computer on June 30, 2017 = Cost of the asset - Depreciation till date =  $39300 - $27510 = $11790

Sale proceeds of the computer = $15400

Gain on Sale of computer = Sale value - Written down Value = $15400-$11790 = $3610

3.  Discarded a delivery truck that was purchased on January 1, 2013. The truck cost $42,900. It was depreciated based on a 6-year useful life with a $3,000 salvage value.

Date Account Name Debit ($) Credit ($)
Dec 31, 2017 Accumulated Depreciation $26600
Loss on asset disposal $16300
Truck $42900
(Being Asset disposed off and loss recorded)

Working:

Depreciation on Truck from (Jan 1, 2013 to Dec 31, 2017) i.e 4 years= Cost of the asset - Salvage Value / Life of the asset = $42900-$3000/6= $6650 * 4 years = $26600

Written down value of the Truck on Dec 31, 2017 = Cost of the asset - Depreciation till date =  $42900 - $26600 = $16300